### v3

# Critical re-examination of a claimed challenge of Bohmian mechanics

Abstract
A recent experiment by Sharoglazova et al. was interpreted by the authors as a challenge to the basic idea of Bohmian mechanics, the phase-velocity relation $\vec{v}(\vec{r},t)=\frac{\hbar}{m}\vec{\nabla}S(\vec{r},t)$. Here, we reinterpret their experiment, highlighting the role of the transient regime before the onset of the stationary state. In this way, we show that the experimental data can naturally fit the Bohmian framework when the energy in the evanescent regime is identified with the quantum potential. It is, however, also possible to interpret the same energy as a kinetic energy, in the framework of Nelson’s stochastic quantum mechanics. In conclusion, far from challenging the Bohmian mechanics, we find that the experimental results of Sharoglazova et al. can be interpreted in either ontologies and are therefore not conclusive in selecting one or the other.
I Introduction
Recently, Sharoglazova et al. [1] claimed to have proven a violation of $\vec{v}(\vec{r},t)=\frac{\hbar}{m}\vec{\nabla}S(\vec{r},t)$, the phase-velocity relation of Bohmian mechanics [2], by measuring the speed [3] of an evanescent wave, which is real and therefore must have $\vec{\nabla}S(\vec{r},t)=\vec{0}$ (here, $S(\vec{r},t)$ is the real phase of the wave function $\psi(\vec{r},t)=\rho^{\frac{1}{2}}(\vec{r},t)e^{iS(\vec{r},t)}$).
The main idea at the basis of this measurement is explained in [4]: consider the composed waveguide used in the experiment, schematically depicted in Fig. 1. In the evanescent regime ($E<V_{0}$, with $V_{0}$ the potential step for $x\geq 0$), the wave in $w_{1}$ from $x<0$ penetrates the step at $x=0$ around $y=+a$ as an exponentially decaying wave for $x>0$ (see also Fig. 2). At the same time, it is experimentally found to fill in also the second waveguide ($w_{2}$) with a profile growing with $x$. This, naturally leads to the question: if some density enters the $x>0$ region in the first waveguide without any apparent velocity (in the evanescent state $\rightarrow\vec{\nabla}S=\vec{0}$), and then a density profile is measured in the second waveguide, indicating a spectral-wave displacement from $w_{1}$ to $w_{2}$, can we conclude that there is a velocity associated to this process? and how to combine this velocity to $\vec{\nabla}S=\vec{0}$? Interestingly, [1, 4] have also been able to quantify the population-transfer dynamics through a well-defined expression for the speed (see Section III), giving, in their words [4], ’a clear physical meaning to the notion of a speed for classically forbidden motion’.
Figure 1: Schematic view of the experimental setup of [1]. The beam comes from the negative-$x$ part of $w_{1}$ and finds a potential step $V_{0}$ for $x\geq 0$, as described by the Hamiltonian of Eq. (5). The coupling of the two waveguides for $x>0$ leads to spectral-density motion from $w_{1}$ to $w_{2}$ (see Fig. 2 for the three measured energy regimes).
While the above argument may appear correct, actually it is not, for a subtle reason, that will be quantified in Section III, and that is here summarized as follows: the formation of a density profile at time $t$ in any waveguide is not necessarily the consequence of a current flow at that time $t$. In fact, the onset of the evanescent density profile measured in stationary conditions is the consequence of a current-flow that only existed during the transient regime and disappears in the stationary state, when the current flow equilibrates in the two directions, freezing the density profile created during the transient regime. So, the profile used to deduce the speed of the particle was built by the current in the transient state.
Yet, though this argument explains how a real displacement (thereby, velocity) of particle density is present in transient conditions, this is not the whole the story. In fact, as we will show in Section IV, the interpretation of what is measured in stationary conditions depends on the ontology chosen: in the context of Nelson’s stochastic mechanics it is actually possible to introduce a speed, that naturally fits the experimental findings of [1]. However, far from ’challenging’ Bohmian mechanics, the explanation in terms of Nelson’s antidiffusive (non-classical!) speed, constitutes a bridge towards the explanation of the experiment [1] also in terms of Bohmian mechanics, as Nelson’s antidiffusion energy is equivalent to Bohmian quantum potential. It is the aim of Section IV to interpret the experiment of [1] in the framework of the two above interpretations and also of the orthodox theory (we remind that, because of the mathematical identity $\vec{\nabla}S=\frac{m}{\hbar\rho}\vec{j}$, if there is no flow in the Bohm interpretation of quantum mechanics, the same is true for the orthodox interpretation).
The paper is organized as follows: Section II is devoted to a one-dimensional introduction to the transient phenomena that will later be used for the full two-dimensional Hamiltonian (5). Though this part is well known in the literature [5, 6, 7], we are not sure that it is also widely known: indeed, the conceptual ideas to explain how transient conditions allow obtaining the evanescent density are already contained in this one-dimensional example and it is therefore worthwhile to linger on it. In Section III, we describe both the stationary and transient solutions of the full two-dimensional hamiltonian (Eq. (5)) and show how to explain the experimental findings of [1]. We underline that the full two-dimensional Schrödinger equation is necessary to describe transversal motion in the $y$ direction (it avoids the introduction of exotic interpretations [8]), even though the effective two-channel one-dimensional analysis performed in [4] is sufficeint to explain the longitudinal motion in the $x$ direction. Section IV is dedicated to the analysis of three different onthologies of quantum mechanics (the orthodox, Bohm’s and Nelson’s intepretations), showing why the experiment can be explained in three different ways, and why each explanation is coherent (only) within the corresponding interpretation. This also mean that we cannot use the ontology of an interpretation (e.g., the diffusive, non-classical speed of Nelson’s interpretation) to deduce the failure of another interpretation (e.g., Bohmian, that would instead explain the same experiment in terms of the quantum potential). Finally, in Section V we draw our conclusions highlighting the pedagogical potentials of this experimental setup.
II Detailed analysis of the one-dimensional problem
Before moving to the full bidimensional Schrödinger equation (5), it might be useful here to remember the main outcomes of the simpler one-dimensional case of a step potential, in transient conditions, along the description of, e.g., [5, 6]. Consider the one-dimensional Schrödinger equation in a step potential:
$i\hbar\frac{\partial\psi}{\partial t}=\left[-\frac{\hbar^{2}}{2m}\frac{\partial^{2}}{\partial x^{2}}+V_{0}\theta(x)\right]\psi$ (1)
where $V_{0}$ is a positive constant and $\theta(x)$ the Heaviside function (1 if $x>0$ and 0 otherwise). Define as usual $k=\frac{\sqrt{2mE}}{\hbar}$ and $\kappa=\frac{\sqrt{2m(V_{0}-E)}}{\hbar}$. Then, take a rectangular wave packet, with $E<V_{0}$, at the initial time $t=0$: $\psi(x,0)=e^{ik_{0}x}\left(\theta(x+L)-\theta(x)\right)$, with $L\gg 1/k_{0}$. This wave-packet is not strictly monochromatic at $k_{0}$ because of the cutting edge at $x=0$ (the one at x=Lx=-L is not actually necessary in our case, as we are interested only in the transient phenomena at the raising edge: it will be eliminated below by letting $L\rightarrow\infty$). The packet travels towards the positive $x$ direction and reaches the step potential, at $x=0$, in the time interval $t\in[0,T]$, with $T=\frac{L}{v_{g}}$ (and $v_{g}$ the group velocity).
The solution of $\psi(x,t)$ for $x>0$, $t>0$ can be obtained using the retarded Green function $G_{t}(x,x^{\prime};t)$ by standard techniques [5, 6, 7], as detailed below. Start from the stationary, energy-dependent, Green function $G_{E}(x,x^{\prime};E)=\frac{m}{i\hbar^{2}k}T(k)e^{-ikx^{\prime}}e^{-\kappa(k)x}$, valid for $x^{\prime}<0$ and $x>0$. Here $T(k)=\frac{2k}{k+i\kappa(k)}$is the transmission coefficient and $\kappa(k)=\sqrt{\frac{2mV_{0}}{\hbar^{2}}-k^{2}}$. Then Fourier-transform to $G_{t}(x,x^{\prime};t)=\int_{-\infty}^{+\infty}G_{E}(x,x^{\prime};E)e^{-iEt/\hbar}\frac{dE}{2\pi}$. Transform the $E$-integral in a $k$-integral by $E=\frac{\hbar^{2}k^{2}}{2m}$, $dE=\hbar^{2}kdk/m$ and approximate $T(k)\simeq T(k_{0})$ (slowly varying around the pole $k=k_{0}$). In order to perform the integral, we should also approximate [9] $\kappa(k)\simeq\kappa_{0}+\frac{k_{0}}{\kappa_{0}}(k-k_{0})$, where $\kappa_{0}=\kappa(k_{0})$. Then the integral can be formally calculated and we find:
$\!\!\!\!\!\!\psi(x,t)$ $=\int_{-\infty}^{+\infty}dx^{\prime}\psi(x^{\prime},0)G_{t}(x,x^{\prime};t)$ (2)
$=\int_{-L}^{0}dx^{\prime}e^{ik_{0}x^{\prime}}G_{t}(x,x^{\prime};t)$
$\simeq\frac{1}{2}T(k_{0})e^{-\kappa_{0}x}e^{-iE_{0}t/\hbar}{\rm erfc}\left(z(x,t)\right)$
with $z(x,t)\!=\!-\frac{k_{0}}{\kappa_{0}}\sqrt{\frac{m}{2\hbar t}}\left(x-iv_{0}t\right)e^{i\pi/4}$ and $v_{0}=\frac{\hbar\kappa_{0}}{m}$. We remind that $\frac{1}{2}{\rm erfc}\left(z(x,t)\right)\xrightarrow[t\rightarrow+\infty]{}1$, so that we recover in this limit the usual stationary solution $\psi_{\rm st}(x,t)=T(k_{0})e^{-\kappa_{0}x}e^{-iE_{0}t/\hbar}$, characterized by a null spatial-dependent phase ($\partial_{x}S=0$) and no current. Yet, this is not true during the transient period and in particular for $t\rightarrow 0^{+}$, when $z(x,t)\rightarrow-\sqrt{\frac{m}{2\hbar t}}\frac{k_{0}}{\kappa_{0}}xe^{i\pi/4}$.
If we keep only the leading term of the asymptotic expansion of erfc(z): ${\rm erfc}(z)\xrightarrow[|z|\rightarrow+\infty]{}\frac{e^{-z^{2}}}{\sqrt{\pi}z}+o(|z|^{-1})$, we obtain the time-dependent probability density (remark that $z^{2}$ is purely imaginary and equal to 1 in the $t=0^{+}$ limit):
$|\psi(x,t=0^{+})|^{2}\simeq\frac{1}{\pi}|T(k_{0})|^{2}e^{-2\kappa_{0}x}\frac{\kappa_{0}}{k_{0}}\frac{\hbar t}{2mx^{2}}$ (3)
Deriving Eq. (3) with respect to $t$ and integrating the result with respect to $x$, the continuity equation leads to the following expression for the current:
$j(x,t=0^{+})\simeq\frac{1}{\pi}|T(k_{0})|^{2}e^{-2\kappa_{0}x}\frac{\kappa_{0}}{k_{0}}\frac{\hbar}{2mx}$ (4)
We remark (see [5], Fig. 20), that the velocity associated to the above current is not a measurement of a propagation, i.e., the density probability does not grow ’from left to right’ in $x$, as we would have had if the argument of the error function was $x-v_{0}t$ (it would have been the case for $E>V_{0}$)). Instead, the imaginary unit in the expression $x-iv_{0}t$ is the signal of a spatially-weighted diffusion (’from bottom to top’) of several high-frequency components of the Fourier spectrum of the initial wave-function, as we can see from Eq. (3). This is the consequence of the parabolic character of the Schrödinger equation, leading to the instantaneous diffusion (with infinite velocity) of all frequency components of the initial wave-packet and a nonzero current with a divergence at $x=0$. This deficiency can be cured describing the system with the relativistic Klein-Gordan equation, allowing a finite propagation speed through a $\theta(ct-x)$ in the Green function (as, for $c\rightarrow\infty$, $\theta(ct-x)\rightarrow 1$, one recovers the nonrelativistic insantaneous increase of the density throughout the whole $x$-axis). Moreover, had we chosen a different profile (e.g. gaussian) for the initial wavefunction, equations (3) and (4) would have been different.
Regardless, the main message (that appears with both the Schrödinger and Klein-Gordan equations, and for any finite-size wavepacket) is that, during the transient period before the stationary state, there is a nonzero density current that builds up the probability density of the evanescent wave. Once the stationary state sets in, the wave-function becomes real and the probability current becomes identically zero. The evanescent density profile at $x>0$ during the stationary state is therefore the result of the above nonzero transient current and of a mechanism allowing to sustain the gradient profile in the subsequent stationary state. As we shall see in Section IV, the nature of this mechanism depends on the chosen ’ontology’ (it is the quantum potential in Bohmian mechanics, the diffusive-speed flow in Nelson’s mechanics, the speed fluctuations in the orthodox mechanics), and cannot be decided through the experimental setup developed in [1].
Before analyzing this part, however, we should repeat the above transient calculations for the realistic bidimensional hamiltonian.
III Analysis of the experiment of [1]
In this section, we analyze the stationary states of the full Schrödinger equation, then analyze the three regimes, depending on the incoming energy, and finally move to the transient states.
III.1 The stationary states of the bidimensional Schrödinger equation
The experimental setup of [1] consisted in a single waveguide in the $x$ direction for $x<0$, that doubled in the $y$ direction for $x>0$ as schematically shown in Fig. 1. The two waveguides are centered at $y=a$ ($w_{1}$) and at y=ay=-a ($w_{2}$). In correspondance to the doubling of the waveguide, for $x\geq 0$, a potential step $V_{0}$ is also added to both $w_{1}$ and $w_{2}$. This experimental setup can be modeled by the Schrödinger equation $i\hbar\partial_{t}\psi(x,y,t)={\hat{H}}(x,y)\psi(x,y,t)$ with the following time-independent Hamiltonian in the $x$ and $y$ directions:
$\!\!\!\!\!\!{\hat{H}}(x,y)$ $=-\frac{\hbar^{2}}{2m}\left(\frac{\partial^{2}}{\partial_{x}^{2}}+\frac{\partial^{2}}{\partial_{y}^{2}}\right)+\theta(-x)\frac{1}{2}m\omega_{0}^{2}(y-a)^{2}$
$+\theta(x)\left[\frac{1}{2}m\omega_{0}^{2}(|y|-a)^{2}+V_{0}\right]$ (5)
where $V_{0}$ is a positive constant representing the step potential at the origin, $\theta(x)$ is the Heaviside function (1 for $x>0$ and 0 for $x<0$), $m$ is the mass of the particle and both waveguides have been modeled in the $y$ direction through harmonic potentials centered at $y=\pm a$. In what follows, we shall only consider the ground states of the harmonic potentials. The stationary eigenstates of Hamiltonian (5) have the form $\psi(x,y)=\theta(-x)\psi_{<}(x,y)+\theta(x)\psi_{>}(x,y)$.
For $x<0$ the longitudinal eigenstates are the plane waves and the transverse eigenstate is the ground-state of the harmonic oscillator ($\chi_{0}(y)=\tfrac{1}{\sqrt{\sigma\sqrt{\pi}}}e^{-\tfrac{y^{2}}{2\sigma^{2}}}$, with $\sigma^{2}=\frac{\hbar}{m\omega_{0}}$):
$\psi_{<}(x,y)=(e^{ik_{0}x}+Re^{-ik_{0}x})\chi_{0}(y-a)$
The total energy is $E=\frac{\hbar\omega_{0}}{2}+\frac{\hbar^{2}k_{0}^{2}}{2m}$, so that the real wavevector $k_{0}=\sqrt{\frac{2m\left(E-\frac{\hbar\omega_{0}}{2}\right)}{\hbar^{2}}}$. To compare with [1, 4], we define $E_{k_{0}}=E-\frac{\hbar\omega_{0}}{2}=\frac{\hbar^{2}k_{0}^{2}}{2m}$, corresponding to their total energy. In the reminder of the paper, we shall remove the constant $\frac{\hbar\omega_{0}}{2}$ term from the counting of the energy, as it appears for both $x<0$ and $x>0$.
For $x>0$, the transverse spectrum consists of a nearly degenerate pair of a symmetric and an antisymmetric states [10] with respect to exchange in $y\pm a$, separated by a tunnel splitting $\hbar\omega_{s}\simeq 2\hbar\omega_{0}\frac{a}{\sigma}e^{-a^{2}/\sigma^{2}}\ll\hbar\omega_{0}$, exponentially small for large $a$ [10]. The symmetric state $\chi_{+}(y)=\frac{\chi_{0}(y-a)+\chi_{0}(y+a)}{\sqrt{2(1+\Omega_{S})}}$ with energy $\varepsilon_{+}=V_{0}-\frac{\hbar\omega_{s}}{2}$ is the ground state, whereas $\chi_{-}(y)=\frac{\chi_{0}(y-a)-\chi_{0}(y+a)}{\sqrt{2(1-\Omega_{S})}}$ has energy $\varepsilon_{-}=\varepsilon_{+}+\hbar\omega_{s}$. Here $\Omega_{S}=e^{-a^{2}/\sigma^{2}}$ is the overlap between $\chi_{0}(y-a)$ and $\chi_{0}(y+a)$. We have kept the overlap explicitly, though it does not play any direct role in the density measurements of [1], because it is necessary in order to have a $j_{y}$ current in the transient state (see Sec. III.B).
The global solution for $x>0$ is:
$\psi_{>}(x,y)=T_{+}\chi_{+}(y)e^{ik_{+}x}+T_{-}\chi_{-}(y)e^{ik_{-}x}$ (6)
where $k_{\pm}=\sqrt{\frac{2m}{\hbar^{2}}(E_{k_{0}}-V_{0}\pm\frac{\hbar\omega_{s}}{2})}$. By imposing the continuity at $(x=0,y=a)$ of the function ($\psi_{<}(0,a)=\psi_{>}(0,a)$) and of the derivatives, we obtain the reflection and transmission coefficients: $R=\frac{k_{0}-\overline{k}}{k_{0}+\overline{k}}$ and $T_{\pm}=\frac{2k_{0}}{k_{0}+\overline{k}}c_{\pm}=Tc_{\pm}$, with $\overline{k}=c_{+}^{2}k_{+}+c_{-}^{2}k_{-}$ and $c_{\pm}=\sqrt{\frac{1\pm\Omega_{S}}{2}}$.
If we put $k_{+}=k_{2}+k_{1}$ and $k_{-}=k_{2}-k_{1}$ (so that $k_{2}=\tfrac{1}{2}(k_{+}+k_{-})$ and $k_{1}=\tfrac{1}{2}(k_{+}-k_{-})$), we can write $\psi_{>}(x,y)$ as:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\psi_{>}(x,y)\!=Te^{ik_{2}x}\left[\cos(k_{1}x)\chi_{0}(y-a)\!+i\sin(k_{1}x)\chi_{0}(y+a)\right]$ (7)
This solution coincides with the solution of [1, 4], if we identify the two (abstract) degrees of freedom $\uparrow$ and $\downarrow$ associated to $\psi(x)$ in [1, 4] with the transversal wavefunctions $\chi_{0}(y-a)$ and $\chi_{0}(y+a)$, respectively (within a phase factor). We remind that $\chi_{0}(y-a)$ and $\chi_{0}(y+a)$ are only approximatively orthogonal (because of the $\Omega_{S}$ overlap), but to a good approximation Eq. (7) reproduces the results of [1, 4]. An equivalent way to look at this equivalence is to construct directly an effective hamiltonian from Eq. (5), by introducing, in the $x>0$ region, the transverse eigenbasis $|\chi_{+}(y)\rangle$ and $|\chi_{-}(y)\rangle$, where the one-dimensional longitudinal motion in the $x$ direction is associated to an abstract two-dimensional space spanned by the eigenbasis $|\chi_{+}(y)\rangle$, $|\chi_{-}(y)\rangle$ representing a linear combination of the two waveguides:
${\hat{H}}_{>}=-\frac{\hbar^{2}}{2m}\frac{\partial^{2}}{\partial x^{2}}+V_{0}+\varepsilon_{+}+\hbar\omega_{s}|\chi_{-}\rangle\langle\chi_{-}|$ (8)
By interpreting $\chi_{0}(y-a)\rightarrow\psi_{\uparrow}$ and $\chi_{0}(y+a)\rightarrow\psi_{\downarrow}$, we recover the full solution of [1, 4] (within an ininfluent phase change and the exact orthogonality that is reached only the limit $\tfrac{a}{\sigma}\gg 1$, when, however, $\omega_{s}\rightarrow 0$). We remark that, though the effective one-dimensional model [4] is simpler for the calculations, it may hide that the intercoupling of the two waveguides corresponds to a transverse ($y$) motion and not to an abstract degree of freedom. Such a transverse motion show up in the transient state, where $j_{y}\neq 0$, as shown below.
We have three different behaviours according to whether: a) $E_{k_{0}}>V_{0}+\frac{\hbar\omega_{s}}{2}$ (Region 1); b) $V_{0}-\frac{\hbar\omega_{s}}{2}<E_{k_{0}}<V_{0}+\frac{\hbar\omega_{s}}{2}$ (Region 2); c) $E_{k_{0}}<V_{0}-\frac{\hbar\omega_{s}}{2}$ (Region 3). In order to compare with [1, 4], we remind that our $\frac{\hbar\omega_{s}}{2}$ corresponds to their $\hbar J_{0}$, and that their $V_{0}$ corresponds to the level where pure propagation starts, i.e., our $V_{0}+\frac{\hbar\omega_{s}}{2}$. Finally, we define $\Delta=E_{k_{0}}-V_{0}$ (corresponding to the $\Delta$ defined in [4]). The full analogy is proposed in Fig. 2.
Figure 2: [Left] Comparison of the energy levels in our notation and in that of [1]. [Right] Two of the three allowed regimes described in the text. For $E_{k_{0}}>V_{0}+\frac{\hbar\omega_{s}}{2}$ (region 1 - upper part), both $k_{+}$ and $k_{-}$ are real and $j_{x}\neq 0$. The beating of the two wavevectors is responsible of the shape of the measured density in $w_{1}$ and $w_{2}$. For $E_{k_{0}}<V_{0}-\frac{\hbar\omega_{s}}{2}$ (region 3 - bottom diagram), both channels become evanescent. It is the most interesting part as the measurement of a speed in this regime is at the basis of the ’challenging’ claim of [1].
Region 1 In this case, both wavevectors are real: $k_{\pm}=\sqrt{\tfrac{2m}{\hbar^{2}}(\Delta\pm\hbar J_{0})}$ and therefore $k_{2}=\frac{\sqrt{m}}{\hbar}\sqrt{\Delta+\sqrt{\Delta^{2}-\hbar^{2}J_{0}^{2}}}$ and $k_{1}=\frac{\sqrt{m}}{\hbar}\sqrt{\Delta-\sqrt{\Delta^{2}-\hbar^{2}J_{0}^{2}}}$, with $\tfrac{\hbar^{2}k_{1}^{2}}{2m}+\tfrac{\hbar^{2}k_{2}^{2}}{2m}=\Delta$ and $k_{1}=\frac{mJ_{0}}{\hbar k_{2}}$, as in [4]. The latter two relations are valid in all three regions.
A current density can be associated to Eq. (7) as: $\vec{j}=\tfrac{\hbar\rho}{m}\vec{\nabla}S(x,y)$ and therefore a speed $\vec{v}=\tfrac{\vec{j}}{\rho}$. The longitudinal current is $j_{x}(x,y)=|T|^{2}\tfrac{\hbar k_{2}}{m}[\cos^{2}(k_{1}x)\chi_{0}^{2}(y-a)+\sin^{2}(k_{1}x)\chi_{0}^{2}(y+a)]+|T|^{2}\tfrac{\hbar k_{1}}{m}\chi_{0}(y-a)\chi_{0}(y+a)$, with the latter term negligible with respect to the former (proportional to the overlap of the wavefunctions in the two waveguides). So the longitudinal current is basically characterized by a transmission at the average wavevector $k_{2}=\frac{k_{+}+k_{-}}{2}$, with beatings at the inverse wavelength $k_{1}$. These beatings are out of phase by $\frac{\pi}{2}$ in the two waveguides, so that the $x$-dependence in the auxiliary waveguide $w_{2}$ is $\propto\sin^{2}(k_{1}x)$. In the limit $\tfrac{a}{\sigma}\rightarrow\infty$, there would be no overlap of the two waveguides, so $\omega_{s}\rightarrow 0$ ($J_{0}\rightarrow 0$ in the notation of [1]) and the current in the second waveguide would be zero ($k_{1}\rightarrow 0$).
The transverse component is only proportional to the overlap of the wavefunctions in the two waveguides, as it should (no overlap implies independence of $w_{1}$ and $w_{2}$, and $\chi_{0}(y-a)$ and $\chi_{0}(y+a)$ would be separately eigenstates of the global hamiltonian): $j_{y}(x,y)=-|T|^{2}\tfrac{\hbar a}{m\sigma^{2}}\sin(2k_{1}x)\chi_{0}(y-a)\chi_{0}(y+a)$ (the negative sign shows a transfer from $w_{1}$ to $w_{2}$ for small values of $x$). We remark that, though $\partial_{x}j_{x}\neq 0$ and $\partial_{y}j_{y}\neq 0$, the current $\vec{j}$ is divergenceless (as $\partial_{x}j_{x}=-\partial_{y}j_{y}$) and therefore it cannot displace spectral weight, as obvious in stationary conditions.
Region 2 As above, $k_{+}$ is real but $k_{-}=iq_{-}$, with $q_{-}=\sqrt{\tfrac{2m}{\hbar^{2}}(-\Delta+\hbar J_{0})}$ real. This implies that $k_{2}=\frac{\sqrt{m}}{\hbar}\sqrt{\Delta+i\sqrt{-\Delta^{2}+\hbar^{2}J_{0}^{2}}}=k_{1}^{*}$. The wavefunction is given by: $\psi_{>}(x,y)=T\allowbreak\left[\chi_{0}(y-a)(e^{i(k_{+}x}+e^{-q_{-}x})+\chi_{0}(y+a)(e^{i(k_{+}x}-e^{-q_{-}x})\right]$. The current in the longitudinal direction is: $j_{x}(x,y)=\tfrac{\hbar}{m}|T|^{2}\big[k_{+}c_{+}^{2}\chi_{+}^{2}+e^{-q_{-}x}(k_{+}\cos(k_{+}x)+q_{-}\sin(k_{+}x))\allowbreak(\chi_{0}^{2}(y-a)-\chi_{0}^{2}(y+a))/4\big]$. The transversal current is: $j_{y}=\tfrac{\hbar a}{m\sigma^{2}}|T|^{2}e^{-q_{-}x}\sin(k_{+}x)\chi_{0}(y-a)\chi_{0}(y+a)$. We remark that $j_{y}$ only depends on the overlap of $\chi_{0}(y-a)$ and $\chi_{0}(y+a)$. Again, as in the previous case, though $\partial_{x}j_{x}\neq 0$ and $\partial_{y}j_{y}\neq 0$, the current density $\vec{j}$ is divergenceless.
Region 3 This is the critical region, as it ’challenges’ Bohmian mechanics. In this case, both wavevectors $k_{+}=iq_{+}$ and $k_{-}=iq_{-}$ are imaginary and there is no stationary current. We have $q_{\pm}=\sqrt{\tfrac{2m}{\hbar^{2}}(-\Delta\mp\hbar J_{0})}$, with $q_{-}>q_{+}>0$. We obtain therefore: $q_{1,2}=\frac{\sqrt{m}}{\hbar}\sqrt{-\Delta\mp\sqrt{\Delta^{2}-\hbar^{2}J_{0}^{2}}}$ (- for $q_{1}$ and ++ for $q_{2}$). We remark that, when $|\Delta|\gg\hbar J_{0}$, $q_{1}\rightarrow 0$ and $q_{2}\rightarrow\tfrac{2m}{\hbar^{2}}|\Delta|$. The wave function in this case is:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\psi_{>}$ $(x,y)=T\left(c_{+}\chi_{+}(y)e^{-q_{+}x}+c_{-}\chi_{-}(y)e^{-q_{-}x}\right)$ (9)
$\!\!\!\!\!\!\!\!=Te^{-q_{2}x}\left[\cosh(q_{1}x)\chi_{0}(y-a)+\sinh(q_{1}x)\chi_{0}(y+a)\right]$
Density measurement in the three regions: We remind the operational measurement of the velocity, for all three regimes, performed in [1], through a relative population analysis: the relative density of the waveguide $w_{2}$ is compared to the full measured density as:
$\rho_{a}(x)=\frac{|\psi_{>}(x,-a)|^{2}}{|\psi_{>}(x,a)|^{2}+|\psi_{>}(x,-a)|^{2}}$ (10)
We can evaluate this quantity for the three regimes, using $\chi_{0}(\pm 2a)\ll\chi_{0}(0)$. Through the above definitions, reminding that $k_{1}=\tfrac{mJ_{0}}{\hbar k_{2}}$ in all three regions, we obtain the same dependences as [4]:
  • Region 1: $\rho_{a}=\sin^{2}(k_{1}x)\rightarrow(k_{1}x)^{2}=\left(\frac{mJ_{0}}{\hbar}\right)^{2}\frac{x^{2}}{k_{2}^{2}}$ for $x\ll 1/k_{1}$.
  • Region 2: $\rho_{a}=\frac{\cosh(q_{-}x)-cos(k_{+}x)}{2\cosh(q_{-}x)}\rightarrow\frac{q_{-}^{2}+k_{+}^{2}}{4}x^{2}=\frac{mJ_{0}}{\hbar}x^{2}$ for $x\ll 1/q_{-},1/k_{+}$. It can be rewritten as: $\rho_{a}=\left(\frac{mJ_{0}}{\hbar}\right)^{2}\frac{x^{2}}{k_{J}^{2}}$, with $k_{J}=\sqrt{\frac{mJ_{0}}{\hbar}}$, independent of $\Delta$.
  • Region 3: $\rho_{a}=\sinh^{2}(q_{1}x)\rightarrow(q_{1}x)^{2}=\left(\frac{mJ_{0}}{\hbar}\right)^{2}\frac{x^{2}}{q_{2}^{2}}$ for $x\ll 1/q_{1}$.
Rewriting $k_{2}$ and $q_{2}$ in terms of $\Delta$ in regions 1 and 3, we have: $\rho_{a}=\frac{mJ_{0}^{2}x^{2}}{|\Delta|(1+\sqrt{1-\alpha^{2}})}$, with $\alpha=\tfrac{\hbar J_{0}}{|\Delta|}$, for both regions. The dependence of $k_{2}$ (region 1), $k_{J}$ (region 2) and $q_{2}$ (region 3) in terms of $\Delta$ is reproduced in Fig. 3. It perfectly fits the experimental data without any need to invoke a speed: the reason is that the measurements of [1] in a stationary regime are primarily measurements of a stationary wavevector. They might be interpreted in terms of an associated speed, but they might even be not (see Section IV). We remark that, in order to force a square-root dependence between the speed and $|\Delta|$ for all values of $|\Delta|$, the experimental data in [1] were modeled by $v=\sqrt{\tfrac{2|\Delta|}{m}}$ (see blue line of Fig. 3), even though the best model is clearly the one obtained here (and in [4]): $k_{2}=\tfrac{\sqrt{m|\Delta|}}{\hbar}$, if $\tfrac{|\Delta|}{\hbar J_{0}}\geq 1$ and $k_{2}=\sqrt{\frac{mJ_{0}}{\hbar}}$, if $\tfrac{|\Delta|}{\hbar J_{0}}\leq 1$ (red line, Fig. 3).
Figure 3: The stationary experimental data (red circles) of [1] can be reproduced in the 3 energy regions using Eq. (10), as a function of $|\Delta|$ (in units $\hbar J_{0}$), both in terms of a stationary wavevector $|k_{2}|$ associated to Bohm’s quantum potential (in the framework of Bohm’s quantum mechanics), left scale of the ordinates, or of a nonclassical speed $|\vec{u}|$, corresponding to Nelson’s antidiffusive speed in his stochastic quantum mechanics, right scale of the ordinates. Here (dashed blue) $|k_{2}|=\sqrt{2m|\Delta|}/\hbar$ if $\tfrac{|\Delta|}{\hbar J_{0}}\geq 1$ and $k_{2}=k_{J}$ if $\tfrac{|\Delta|}{\hbar J_{0}}\leq 1$, whereas $|\vec{u}|=\sqrt{\tfrac{2|\Delta|}{m}}$ if $\tfrac{|\Delta|}{\hbar J_{0}}\geq 1$ and $|\vec{u}|=k_{J}$ if $\tfrac{|\Delta|}{\hbar J_{0}}\leq 1$. The black dots correspond to the square-root model of [1].
The increasing dependence of the wavevector $q_{2}$ as a function of $\Delta$ in the evanescent region 3 is a consequence of the Heisenberg principle: the higher the energy, the more localised the evanescent wavepacket and therefore the higher $q_{2}$. When, as done in [1, 4], a speed is associated to the evanescent wave-vector through the relation $v=\tfrac{\hbar q_{2}}{m}$, this property is herited by the speed, that becomes higher the lower is the energy $E_{k_{0}}$ compared to $V_{0}$ (so, the higher is $|\Delta|$). The question is: does this speed correspond to the motion of anything? We shall answer to this question in Section IV, where we associate the non-classical speed to $q_{2}$ in the framework of Nelson’s quantum mechanics [11]. At the same time, we shall see that it is also possible to explain the experiment in the framework of Bohmian mechanics without introducing any speed at all, by interpreting the wavevectors $q_{1}$ and $q_{2}$ as manifestations of the stationary quantum potential.
But before we have to answer the question raised in the Introduction: how does the spectral weight move from the first to the second waveguide in the evanescent region in the absence of any speed? Does it in stationary conditions? The key hypothesis in [1] is, in their words: ’If the probability amplitude is initially concentrated entirely in one of the states [here, $w_{1}$], the population in the initially unoccupied state [here, $w_{2}$] follows $\sin^{2}(J_{0}t)$ as a function of time’. This is not true in stationary conditions - it is only true in the transient regime, as we shall see it in the next subsection.
III.2 The transient regime for the bidimensional Schrödinger equation
The reason to study the transient regime is simple. From the continuity equation, $\partial_{t}\rho(x,y,t)=-\vec{\nabla}\cdot\vec{j}$, integrating both terms in time from $t=0$ (when there is no density for $x>0$) to $\infty$ (when the stationary state is reached), we get:
$\rho(x,y)=-\int_{0}^{\infty}dt\left(\partial_{x}j_{x}+\partial_{y}j_{y}\right)$ (11)
where $\rho(x,y)$, time independent, is the stationary density. Of course, during this transient regime, the spectral weight has to move transversally from $w_{1}$ (where it enters at $x=0$, $t=0$) to the auxiliary waveguide $w_{2}$, so as to reproduce the stationary density at later times. This can only be done through a nonzero time-integral of $j_{y}$ (whereas its longitudinal distribution is obtained through the combined action of the time-integrals of $j_{x}$ and $j_{y}$ from Eq. (11)). Though the precise time-dependence of the current density requires a precise knowledge of the initial wavepacket, we can show below that, for a simplified form of the initial wavepacket (as in Section II), there is a nonzero $j_{y}$ during the transient regime that, in the stationary limit, reduces to the expression of $\rho_{a}$.
As already done for the one-dimensional case, we first evaluate the energy-dependent Green function, that obeys: $(E-\hat{H})G_{E}(x,x^{\prime};y,y^{\prime})=\delta(x-x^{\prime})\delta(y-y^{\prime})$. As we need to evaluate the transition of a particle initially in the negative-x region and in the main waveguide towards the positive x-region and the auxiliary waveguide, we take $x^{\prime}<0$, $y^{\prime}>0$, $x>0$ and any $y$. The Green function takes the form $G_{E}(x,x^{\prime};y,y^{\prime})=\Sigma_{\alpha=\pm}\chi_{\alpha}(y)\chi_{\alpha}(y^{\prime})g_{\alpha}(x,x^{\prime};E)$, where $g_{\pm}(x,x^{\prime};E)$ is the solution of the one-dimensionnal, channel-dependent, Green function equation: $\left[E_{k_{0}}+\frac{\hbar^{2}}{2m}\partial_{x}^{2}-E_{\pm}\right]g_{\pm}(x,x^{\prime};E)=\delta(x-x^{\prime})$, with $E_{\pm}=V_{0}\mp\frac{\hbar\omega_{s}}{2}$, as $x>0$.
In the evanescent region $E_{k_{0}}<V_{0}-\frac{\hbar\omega_{s}}{2}$, and the solution is: $g_{\pm}(x,x^{\prime};E)=\frac{m}{i\hbar^{2}k}T_{\pm}(k)e^{ikx^{\prime}}e^{-q_{\pm}(k)x}$, like in the one-dimensional case of Section II, except for the channel dependence. Here $T_{\pm}(k)=\frac{2k}{k+i\overline{q}}$, with $\overline{q}=q_{+}c_{+}^{2}+q_{-}c_{-}^{2}$ and $q_{\pm}(k)=\sqrt{\frac{2m}{\hbar^{2}}\left(V_{0}-E_{k_{0}}\mp\frac{\hbar\omega_{s}}{2}\right)}$.
Therefore, the general solution for the wave function is obtained by summing over the two one-dimensional Green functions (one for each transverse eigenchannel), with the same hypothesis already used in Section II (in particular, the linear expansion of $q(k)$). We get:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\psi(x,y,t)=\frac{e^{-\frac{iE_{0}t}{\hbar}}}{2}\!\!\!\sum\limits_{\alpha=\pm}\!c_{\alpha}\chi_{\alpha}(y)T_{\alpha}(k_{0})e^{-q_{\alpha}^{(0)}x}{\rm erfc}(z_{\alpha}(x,t))$ (12)
where $z_{\alpha}(x,t)=-\frac{k_{0}}{q_{\alpha}^{(0)}}\sqrt{\frac{m}{2\hbar t}}\left(x-iv_{0\alpha}t\right)e^{i\pi/4}$, and $v_{0\alpha}=\frac{\hbar q_{\alpha}^{(0)}}{m}$ have the same expression as in Section II for a single channel. As in that case, the total energy $E=E_{k_{0}}+\frac{\hbar\omega_{0}}{2}$ is just an overall phase factor (as it should, in stationary conditions) and does not contribute to the density nor to the current: all the time-dependence comes from the complementary error function ${\rm erfc}(z_{\alpha}(x,t))$. The only difference is the channel dependence $\alpha$ implying cross contributions in the calculation of the total density and of the density current (and in general of all measurable quantities, that have a form $\psi^{*}{\hat{O}}\psi$, for any operator $^{\hat{O}}$).
In particular, we obtain the following expression for $j_{y}(x,y,t)$:
$j_{y}(x,y,t)=\frac{\hbar}{m}\Im[\psi^{*}(x,y,t)\partial_{y}\psi(x,y,t)]$ (13)
$\!\!\!\!\!\!=\frac{-\hbar a}{4m\sigma^{2}}|T|^{2}e^{-\tfrac{q_{2}x}{2}}\chi_{0}^{(-)}\chi_{0}^{(+)}\Im[{\rm erfc}^{*}(z_{+}){\rm erfc}(z_{-})]$
where $\chi_{0}^{(\pm)}=\chi(y\pm a)$. We remark that, as it should, for $t\rightarrow\infty$, the erfc becomes a real function and so $j_{y}$ is zero. Moreover, $j_{y}$ decreases longitudinally as $e^{-q_{2}x/2}$ and is initially negative, so, from $w_{1}$ to $w_{2}$, as $\Im[{\rm erfc}^{*}(z_{+}){\rm erfc}(z_{-})]$ is initially positive. These are the main features of Eq. (13) in this specific case of sudden jump in the pulse: a detailed analysis would require a precise knowledge of the initial wavepacket (here we used the simplyfying rectangular packet), but is finally not required for the main message of this paper.
The conclusion is that the stationary density in both waveguides $w_{1}$ and $w_{2}$ is ’built’ during the transient regime by the nonvanishing current flowing in this regime. Once this is done, one might wonder how a non-constant stationary density can ’freeze’, for example, in the exponential spatial decay of region 3. The explanation of this is different, according to the interpretation of quantum mechanics chosen: in the orthodox interpretation (as usually known)it is an eigenstate of the Schrödinger equation; in Bohm’s interpretation, there is a quantum force, associated to a quantum potential $Q(x,y)$, blocking the particle in the given density distribution; finally, in Nelson’s interpretation, the spatial exponential decay of region 3 is due to a non-classical speed, acting antidiffusively, that forces the particle to distribute in space according the solution of the Schrödinger equation, as we shall see in the next Section.
IV The interpretation in Bohm’s and Nelson’s frameworks.
The average kinetic energy in the orthodox interpretation of quantum mechanics is: $E_{\rm kin}=-\frac{\hbar^{2}}{2m}\int dV\psi^{*}\nabla^{2}\psi=+\frac{\hbar^{2}}{2m}\int dV\vec{\nabla}\psi^{*}\cdot\vec{\nabla}\psi$, where $dV$ stands for volume integration and, after the integration by parts, we have removed a divergence term, as possible in stationary conditions. By replacing $\psi=\sqrt{\rho}e^{iS}$, multiplying and dividing by $\rho=\psi^{*}\psi$ and using the identity $(\nabla\rho)^{2}=\psi^{2}(\nabla\psi^{*})^{2}+(\psi^{*})^{2}(\nabla\psi)^{2}+2\vec{\nabla}\psi^{*}\cdot\vec{\nabla}\psi$, we finally get [14]:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!E_{\rm kin}\!=\!\!\int\!\!dV\rho(\vec{r},t)\!\left[\frac{\hbar^{2}}{2m}(\vec{\nabla}S)^{2}+\frac{\hbar^{2}}{8m}\frac{(\vec{\nabla}\rho)^{2}}{\rho^{2}}\right]\!=E_{\rm kin}^{\rm cl}+Q$ (14)
The first term in Eq. (14) is Bohm’s average kinetic energy (corresponding to the classical kinetic energy, $E_{\rm kin}^{\rm cl}$) and the second is the average energy that Bohm associated to a quantum potential $Q$ in equation (9) of [2]. So, the onthology of Bohm’s theory is the association, to what in the orthodox theory would appear as kinetic-energy fluctuations, with a ’quantum’ potential energy [13]. We can describe the same average kinetic energy with a different onthology in Nelson’s stochastic quantum mechanics [11], where the same term $Q$ is associated to a non-classical speed, $\vec{u}=\tfrac{\hbar}{2m}\vec{\nabla}\ln\rho=+\tfrac{\hbar}{2m}\tfrac{\vec{\nabla}\rho}{\rho}$, where the positive sign in the last expression has been explicitly written to underline the non-classical, antidiffusive, character of this velocity (it goes towards the density gradient instead of opposing to it as in usual diffusion). In the Nelson’s interpretation, the total velocity of the quantum particle is given by the sum of two terms, $\vec{v}_{\rm Tot}=\vec{v}+\vec{u}$. The first term is the center-of-mass velocity $\vec{v}=\frac{\hbar}{m}\vec{\nabla}S(\vec{r},t)$ and the second velocity $\vec{u}$, depending on the amplitude gradient, represents a fluctuating hidden-variable contribution [12], averaging to zero, but leading to the nonzero variance [15]:
$\frac{1}{2}m\langle\vec{u}\cdot\vec{u}\rangle=\int dV\rho(\vec{r},t)\left(\frac{\hbar^{2}}{8m}\frac{(\vec{\nabla}\rho)^{2}}{\rho^{2}}\right)$ (15)
So, the energy associated to the quantum potential $Q$ can be interpreted either as a form of non-classical potential energy, in Bohm’s framework, or as a form of non-classical kinetic energy, in Nelson’s framework. If we evaluate Eq. (14) in the energy regions 1 and 3 (and neglect the quantum potential associated to the transverse motion $Q_{y}$ [16]), we get, in the energy region 1: $E_{\rm kin}^{\rm cl}=\tfrac{\hbar^{2}k_{2}^{2}}{2m}$ and $Q=\tfrac{\hbar^{2}k_{1}^{2}}{2m}$, and, in the energy-region 3: $E_{\rm kin}^{\rm cl}=0$ and $Q=\tfrac{\hbar^{2}q_{1}^{2}}{2m}+\tfrac{\hbar^{2}q_{2}^{2}}{2m}$. So, in the latter case where the wavefunction is real and $S(x,y)=0$, we can identify $\Delta=Q$, i.e., the parameter $\Delta$ of [1] measures the integrated value of the quantum potential. From this, it is clear that, in the energy region 3, we might read Fig. 3 as a speed vs. energy, but also, in Bohm’s interpretation, as a wavevector ($q_{2}$) vs. quantum potential ($Q$) relationship. As such, the experiment [1] is silent on which interpretation (Bohm’s or Nelson’s) we should choose. In fact, if we stick to Bohm theory, the $q_{2}(Q)$ relationship of Fig. 3 expresses the way the quantum potential acts so as to keep the density profile the way it is. There is no speed in stationary conditions for an evanescent wave $\vec{\nabla}S=\vec{0}$ and no need to introduce any: as seen in Sec. IIIB, all spectral-weight displacement takes place in the transient regime. Bohm interpretation is fully consistent also in the energy region 3, by interpreting Fig. 3 as $q_{2}(Q)$.
Of course, it is also possible to prefer instead Nelson’s approach and, in this case, it is indeed possible to introduce a speed $u$ that, in region 3, can be associated to $q_{2}$ through the above relations: $\tfrac{mu^{2}}{2}=\tfrac{\hbar^{2}q_{2}}{2m}$, leading to $u=\tfrac{\hbar q_{2}}{m}$. Nelson’s approach is the ontology of [1, 4]. It is possible, but not necessary. And it does not challenge Bohmian mechanics.
V Discussion and Conclusions
The final picture of the experiment [1] is the following: when the initial wavefront reaches the potential step at $x=0$, a transient regime sets in where the density profile, measured during the subsequent stationary regime, starts forming (e.g., in $w_{2}$). The corresponding transient, time-dependent, wave function is described in Section III.B. During this time there exists a nonzero current density which is not divergenceless and allows the formation of the spectral density ($\partial_{t}\rho(x,y,t)\neq 0$) in the main and auxiliary waveguides ($w_{1}$ and $w_{2}$). The continuity equation (11) shows that the final, normalized, density $\rho_{a}(x)\simeq\rho(x,-a)$ measured in [1] is formed by the time-integral of the opposite of the density-current divergence. In particular, a nonzero $j_{y}(x,y)$ is required to transport spectral weight from $w_{1}$ to $w_{2}$, as shown in section III.B.
Once the transient regime leaves the place to stationary conditions, a divergenceless density current sets in in energy-regions 1 and 2 (i.e. $\partial_{t}\rho(x,y,t)=0$ everywhere). Spectral weight does not change any more and the center-of-mass speed is zero, in any interpretation (orthodox, Nelson’s, Bohm’s): measurable quantities in the stationary regime cannot be associated to non-divergenceless transport. In Nelson’s theoretical framework, it is possible to define a non-classical, divergenceless, speed ($\vec{u}$, Eq. (15)) and interpret in this way Fig. 3, as done in [1]. In Bohm’s theoretical framework, instead, the quantum potential describes the same fact (the parabolic increasing of the density of $w_{2}$ in the $x$-direction) through a modulated (static) wavevector, without the need of a speed.
So, the conceptual misinterpretation in [1] has been to associate a speed at time $t$ to the measured density profile at time $t$. Instead, as we have seen, the invoked displacement from $w_{1}$ to $w_{2}$ at $x=0^{+}$ takes place before, in the transient regime. With this remark, we are no more forced to think that something is moving and we can explain the experiment in stationary conditions within Bohm ontology in terms of the relation between a static wavevector and the quantum potential. As seen, the interpretation in terms of a nonclassical speed proposed by [1] is also valid, but within the ontology proposed in Nelson’s stochastic quantum mechanics.
To conclude, even though the challenge to Bohmian interpretation disappears with the analysis proposed here, we think that, once correctly intepreted, the experimental setup and the measurements of Sharoglazova et al. are pedagogically relevant because they provide clear experimental data on a quantum mechanical problem with unusual geometry (a potential-step in the $x$ direction coupled to a tunnel barrier in the $y$ direction) where it is possible to build a relatively simple model Hamiltonian (Eq. (5)) to analyze the system both in stationary conditions and in the transient regime with almost analytical calculations. As a non-negligeable add-on, it leads to a careful discussion about the meaning of the quantum potential, in Bohm’s interpretation, and of the non-classical diffusion velocity, in Nelson’s interpretation and their relations with the kinetic-energy fluctuations in the orthodox interpretation, so as to give a complete example of quantum mechanics useful to undergraduate and graduate students (and researchers) in the field.